OpenStudy (anonymous):

The minimum and maximum temperature on a cold day in Lollypop Town can be modeled by the equation below: 2|x − 6| + 14 = 38 What are the minimum and maximum temperatures for this day? x = −9, x = 21 x = −6, x = 18 x = 6, x = −18 No solution

OpenStudy (anonymous):

\[2|x-6|+14=38\Rightarrow 2|x-6|=24\Rightarrow |x-6|=12\Rightarrow x-6 = 12 \text{ or } x-6 = -12\]

OpenStudy (anonymous):

solve for x in each case. I assume x is the max or min temp.

OpenStudy (anonymous):

Ohhh , Okay thank you

OpenStudy (anonymous):

you're welcome